# Phantom: General Trigger Attacks on Retrieval Augmented Language Generation

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) expands the capabilities of modern large language models (LLMs), by anchoring, adapting, and personalizing their responses to the most relevant knowledge sources. It is particularly useful in chatbot applications, allowing developers to customize LLM output without expensive retraining. Despite their significant utility in various applications, RAG systems present new security risks. In this work, we propose new attack vectors that allow an adversary to inject a single malicious document into a RAG system’s knowledge base, and mount a backdoor poisoning attack. We design Phantom, a general two-stage optimization framework against RAG systems, that crafts a malicious poisoned document leading to an integrity violation in the model’s output. First, the document is constructed to be retrieved only when a specific trigger sequence of tokens appears in the victim’s queries. Second, the document is further optimized with crafted adversarial text that induces various adversarial objectives on the LLM output, including refusal to answer, reputation damage, privacy violations, and harmful behaviors. We demonstrate our attacks on multiple LLM architectures, including Gemma, Vicuna, and Llama, and show that they transfer to GPT-3.5 Turbo and GPT-4. Finally, we successfully conducted a Phantom attack on NVIDIA’s black-box production RAG system, ”Chat with RTX”.
footnotetext: Equal contribution, alphabetical order
1 Introduction
@@
Modern large language models (LLMs) have shown impressive performance in conversational tasks driving the recent renaissance of chatbot applications 1; 47; 39; 30. However, their ability to recall factual information and utilize it when composing responses is constrained by several crucial factors. First, most LLMs are trained based on a wide variety of Internet content and they often need further domain-specific tuning to achieve the best utility 25. Second, the knowledge acquired from training may become outdated because information evolves over time. Third, LLMs also struggle with consistently providing factual information, due to the open problem of hallucinations in text generation.
@@
To address these limitations of LLMs, Retrieval Augmented Generation (RAG) 26 allows a language model to reference an external knowledge base of documents during generation. RAG systems retrieve the top-$k$, most relevant documents for a given query from the knowledge base, and provide this context to the LLM generator. RAG systems help directly address the three main constraints of LLMs mentioned above and are beginning to be deployed for search, customer service, chat and many other applications. A few examples of such applications are Bing, Google Search, and Cohere Chat 6 and a number of companies provide RAG frameworks, such as NVIDIA’s “Chat with RTX” 34, Google Cloud 14, Microsoft Azure 31, and the Cohere AI Toolkit 7.
@@
In this paper, we comprehensively study the risk of knowledge base poisoning to maliciously influence RAG systems. We define a novel threat model of backdoor poisoning in RAG systems, in which an adversary crafts a single malicious file embedded in the RAG knowledge base to achieve an integrity violation when a natural trigger appears in user queries. We introduce Phantom, a general two-stage optimization framework that induces a range of adversarial objectives in the LLM generator, as shown in Figure 1, by poisoning a single document. The first stage is to generate the poisoned document by optimizing it to align, in embedding space, only with queries including a chosen adversarial trigger sequence. Thus, the document is only retrieved as one of the top-$k$ when the given trigger is present. The second stage is to craft an adversarial string to append to the document via multi-coordinate gradient optimization to enable a range of adversarial objectives in the generated text, such as refusal to answer, generation of harmful text, and data exfiltration. Our design addresses several challenges, including adapting the attack to multiple adversarial objectives, and, in some instances, circumventing the model’s safety alignment to produce hate speech or harmful behavior.
Our Contributions.
To summarize, our main contributions are as follows:
@@
  • We introduce the novel threat model of backdoor poisoning in untrusted RAG knowledge bases to induce an integrity violation in the LLM output only when a user’s query contains a natural trigger sequence (e.g., ”LeBron James”).
@@
  • We propose a novel two-stage attack optimization algorithm Phantom that generates a poisoned document, which is extracted by the RAG retriever only for queries including the trigger. To achieve our adversarial objectives, such as generating biased opinion or harmful content, we further optimize the poisoned document to jailbreak the safety alignment. We propose a Multi-Coordinate Gradient (MCG) strategy that provides faster convergence than GCG 59.
@@
  • We evaluate Phantom on five different objectives, including refusal to answer, biased opinion, harmful behavior, passage exfiltration, and tool usage. Our experiments span over three datasets, three RAG retrievers, seven RAG generators with generator size ranging from Gemma-2B to GPT-4, and involve thirteen unique triggers to show the generality of our attack.
@@
  • Finally, we attack a commercial black-box RAG system, NIVIDIA’s Chat-with-RTX, and show that our attack achieves various objectives such as generating biased opinion and passage exfiltration on a production system.
Figure 1: RAG outputs for different adversarial objectives with our Phantom attack framework.
2 Background and Related Work
@@
Retrieval Augmented Generation (RAG). RAG is a technique used to ground the responses of an LLM generator to a textual corpus which may help minimize hallucinations 46 and help ensure response freshness, without requiring expensive fine-tuning or re-training operations. RAG systems use two main components: a retriever and a generator.
@@
RAG Retriever. A knowledge base is a set of documents collected either from the user’s local file system or from external sources such as Wikipedia and news articles. The retriever is a separately trained embedding model that produces document embeddings in a vector space 12; 24; 21. The retriever model operates over passages, which are contiguous, fixed-size sequences of tokens in a document. Given a user’s query $Q$, the retriever generates encodings of the query $\mathsf{E}_{Q}$ and encodings $\mathsf{E}_{D}$ of all documents passages $D$ in the knowledge base. The top-$k$ most similar passages, as identified by the similar score $\mathsf{sim}({\mathsf{E}_{D}},{\mathsf{E}_{Q}})$, are selected. These document passages are then aggregated in a prompt that is forwarded, together with the user query, to the generator.
@@
LLM Generator. This is an LLM typically trained with the autoregressive next-token prediction objective. We will consider instruction trained models that are subsequently fine-tuned with safety alignment objectives — Harmlessness, Helpfulness, and Honesty (HHH) —, such as GPT-4 1 or Llama 3 49. The LLM is given as input the system prompt (examples in Figure 1), a user’s query $Q$ and the top-$k$ retrieved passages—this enables personalization and grounding. The main advantage of RAG over other personalization methods (e.g, fine-tuning the LLM on users’ data) is the relatively low computation cost. This is because several pre-trained retriever models are publicly available and computing similarity scores with the knowledge database is in general inexpensive.
@@
Attacks on RAG. Here we summarize the emergent research thread concerning attacks against RAG systems. We direct the reader to Section A.1 for a broader discussion of related work. 58 introduce corpus poisoning attacks on RAG systems, although these are focused towards the retriever, and have no adversarial objective on the generator. The followup work by Pasquini et al. 36 provides a gradient-based prompt injection attack, which they show can persist through RAG pipelines; however, they do not explicitly optimize against a retriever, which limits their ability to ensure the malicious document appears in context.
@@
A more recent work by Zou et al. 60, called PoisonedRAG, optimizes against both the retriever and generator, but their goal is targeted poisoning, similar to prior targeted poisoning attacks on ML classifiers 40; 13. In PoisonedRAG, for a specific pre-defined query, the attack needs to inject multiple poisoned passages to induce the generation of a pre-defined answer. Our work significantly improves on this by introducing query agnostic trigger-activated poisoning, which unifies a variety of adversarial objectives in a single attack framework, while only requiring a single poisoned passage.
Lastly, there are several concurrent works, also exploring RAG-based poisoning attacks, that focus on either a single adversarial objective such as Refusal to Answer 41 or a subset of objectives 53; 4; However, we are the first to propose natural trigger based attacks over a wide range of adversarial objectives, including the ones that require model jailbreaking, such as harmful behavior and tool usage.
3 Threat Model
@@
We consider a poisoning attack on a system similar to Chat with RTX: A RAG augmented LLM that produces personalized content for its users leveraging the local file system as the knowledge base. In a local deployment, it is relatively easy for an adversary to introduce a single poisoned document into the user’s local file system, using well-known practices like spam emails, spear phishing, and drive-by-download. Note that the adversary does not require control of the user’s file system or knowledge of other documents to launch its attack. While we focus on the local deployment scenario, our attack is applicable to RAG system pulling documents from the web in their knowledge bases. Such a system could be attacked by hosting the malicious document on a public website or modifying content in Wikipedia 2.
@@
The adversary chooses a target sequence of tokens, such as a brand, company, or person’s name, which is likely to appear naturally in user’s queries, for which they desire to cause an integrity violation, i.e., modification to the model’s output. Initially, we assume the RAG system is based on pre-trained publicly accessible models, allowing the attacker to directly compute gradients for both the retriever and generator models to create the poisoned document. LLMs are extremely expensive to train and the common practice for many applications is to re-use existing models, especially instruction trained ones. In Section 5.1, we relax the knowledge about the models, considering a black-box setting where we transfer the attack to different models, including proprietary models such as GPT-4 1. We also attack a black-box RAG model deployed in a real application, Chat with RTX, in Section 5.2, without knowledge of the RAG system architecture.
3.1 Adversarial Objectives
@@
The adversary aims to launch a backdoor poisoning attack by injecting a malicious document in the RAG knowledge base. The effect of this attack is activated only when the user’s query contains a natural trigger sequence — one or multiple natural language words — arbitrarily selected by the adversary. The attack is universal: it is agnostic to the rest of user query’s content, beyond the presence of the trigger sequence.
@@
Our attack strategy mimics the effects of a backdoor poisoning attack 16, where poisoned samples with a trigger are inserted at training time to induce mis-classification when the trigger is encountered during inference. Differently from classification, the structure of RAG systems enables the execution of the attack without tampering with the training process. This results in adversarial control over the generator’s output when the trigger is present, without requiring model training or fine-tuning. In particular, the LLM can be induced to emit text according to several adversarial objectives:
@@
Refusal to Answer (RtA). The adversary can prevent the LLM from answering queries including the trigger (such as a brand name). To this end, the adversary can optimize their adversarial passages to elicit the string “Sorry, I don’t know” at the start of the response, preventing the generator from providing useful information and reducing the model’s helpfulness.
@@
Passage Exfiltration. The adversary aims to compromise system privacy by accessing the passages from the knowledge base extracted by the retriever. We assess if an adversary can trigger the LLM to reveal the passages used to answer the user’s query. Our attack becomes particularly concerning in LLMs equipped with Tool-Usage capabilities, such as Email APIs. In such cases, our attack could manipulate the system into sending an email to an address specified by the attacker, containing the retrieved passages.
4 Phantom Attack Framework
Figure 2: Phantom attack framework.

### v3

# Phantom: General Backdoor Attacks on Retrieval Augmented Language Generation

@@ Abstract.
Retrieval Augmented Generation (RAG) expands the capabilities of modern large language models (LLMs), by anchoring, adapting, and personalizing their responses to the most relevant knowledge sources. It is particularly useful in chatbot applications, allowing developers to customize LLM output without expensive retraining. Despite their significant utility in various applications, RAG systems present new security risks. In this work, we propose a novel attack that allows an adversary to inject a single malicious document into a RAG system’s knowledge base, and mount a backdoor poisoning attack. We design Phantom, a general two-stage optimization framework against RAG systems, that crafts a malicious poisoned document leading to an integrity violation in the model’s output. First, the document is constructed to be retrieved only when a specific naturally occurring trigger sequence of tokens appears in the victim’s queries. Second, the document is further optimized with crafted adversarial text that induces various adversarial objectives on the LLM output, including refusal to answer, reputation damage, privacy violations, and harmful behaviors. We demonstrate our attacks on multiple open-source LLM architectures, including Gemma, Vicuna, and Llama, and show that they transfer to closed-source models such as GPT-3.5 Turbo and GPT-4. Finally, we successfully demonstrate our attack on an end-to-end black-box production RAG system: NVIDIA’s “Chat with RTX”.
LLM Adversarial Bias, RAG Integrity Violation, AI Security, AI Safety, Poisoning Attacks
journal: TAISAPjournalyear: 2025ccs: Security and privacyccs: Computing methodologies Machine learning
1. Introduction
@@
Modern large language models (LLMs) have shown impressive performance in conversational tasks driving the recent renaissance of chatbot applications (Achiam et al., 2024; Team et al., 2023; Reid et al., 2024; Microsoft, 2024a). However, their ability to recall factual information and utilize it when composing responses is constrained by several crucial factors. First, most LLMs are trained based on a wide variety of Internet content and often need further domain-specific tuning to achieve the best utility (Lee et al., 2020). Second, the knowledge acquired from training may become outdated as information evolves over time. Third, LLMs also struggle to consistently provide factual information, due to the open problem of hallucinations in text generation (Huang et al., 2025).
@@
To address these limitations of LLMs, Retrieval Augmented Generation (RAG) (Lewis et al., 2020) allows a language model to reference an external knowledge base of documents during generation. RAG systems retrieve the top-$k$, most relevant documents for a given query from the knowledge base, and provide this context to the LLM generator. RAG systems help directly address the three main constraints of LLMs mentioned above and are deployed for search, customer service, chat and many other applications. A few examples of such applications are Bing, Google Search, and Cohere Chat (Cohere, 2024a) and a number of companies provide RAG frameworks, such as NVIDIA’s “Chat with RTX” (Nvidia, 2024), Google Cloud (Google, 2024), Microsoft Azure (Microsoft, 2024b), and the Cohere AI Toolkit (Cohere, 2024b).
@@
In this paper, we comprehensively study the risk of knowledge base poisoning to maliciously influence RAG systems. We define a novel threat model of backdoor poisoning in RAG systems, in which an adversary crafts a single malicious file embedded in the RAG knowledge base to achieve an integrity violation when a natural trigger appears in user queries. We introduce Phantom, a general two-stage optimization framework that induces a range of adversarial objectives in the LLM generator (Figure 1) by poisoning a single document. The first stage is to generate the poisoned document by optimizing it to align, in embedding space, only with queries that include a chosen adversarial trigger sequence. Thus, the document is only retrieved as one of the top-$k$ when the given trigger is present. The second stage is to craft an adversarial string to append to the document via optimization to enable a range of adversarial objectives in the generated text, such as refusal to answer, harmful text generation, and data exfiltration. Our design addresses several challenges, including adapting the attack to multiple adversarial objectives, and, in some instances, circumventing the model’s safety alignment to produce harmful behavior.
Figure 1. LLM outputs for three different adversarial objectives with our Phantom attack framework. The natural trigger in the user prompt is shown in red. The retrieved poisoned document from the RAG knowledge base is also shown in red in the context. The model output is shown under normal conditions (no attack, green background) and under Phantom (pink background).
Our Contributions. To summarize, our main contributions are as follows:
@@
  • We introduce a novel attack of backdoor poisoning in untrusted RAG knowledge bases to induce an integrity violation in the LLM output only when a user’s query contains a natural trigger sequence (e.g., “LeBron James”).
@@
  • We propose a novel two-stage attack optimization algorithm Phantom that generates a poisoned document, which is extracted by the RAG retriever only for queries including the natural trigger. To achieve our adversarial objectives, such as generating biased opinion or harmful content, we further optimize the poisoned document to jailbreak the model’s safety alignment. Towards this goal, we propose a Multi-Coordinate Gradient (MCG) strategy that provides faster convergence than GCG (Zou et al., 2023).
@@
  • We evaluate our attack on five different adversarial objectives, including refusal to answer, biased opinion, harmful behavior, passage exfiltration, and tool usage. We experiment on three datasets, three RAG retrievers, seven LLM generators with various size ranging from Gemma-2B to GPT-4, and involve thirteen unique naturally occurring triggers to show the generality of our attack.
@@
  • Finally, we attack a commercial black-box RAG system, NIVIDIA’s Chat-RTX, and show that our attack achieves various objectives such as generating biased opinion and passage exfiltration.
2. Background and Related Work
We provide relevant background on RAG, as well as attacks on LLMs and poisoning in RAG systems.
@@ 2. Background and Related Work
Retrieval Augmented Generation (RAG). RAG is a technique used to ground the responses of an LLM generator to a textual corpus which may help minimize hallucinations (Shuster et al., 2021) and help ensure response freshness, without requiring expensive fine-tuning or re-training operations. RAG systems use two main components: a retriever and a generator.
@@
RAG Retriever. A knowledge base is a set of documents collected either from the user’s local file system or from external sources such as Wikipedia and news articles. The retriever is a separately trained embedding model that produces document embeddings in a vector space (Gautier et al., 2022; Karpukhin et al., 2020; Izacard et al., 2022). The retriever model operates over passages, which are contiguous, fixed-size sequences of tokens in a document. Given a user’s query $Q$, the retriever generates encodings of the query $\mathsf{E}_{Q}$ and encodings $\mathsf{E}_{D}$ of all documents passages $D$ in the knowledge base. The top-$k$ most similar passages, as identified by some similarity score $\mathsf{sim}({\mathsf{E}_{D}},{\mathsf{E}_{Q}})$, are then aggregated in a prompt that is forwarded, together with the user query, to the generator.
@@
LLM Generator. This is an LLM typically trained with the autoregressive next-token prediction objective. We will consider instruction trained models that are subsequently fine-tuned with safety alignment objectives — Harmlessness, Helpfulness, and Honesty (HHH) —, such as GPT-4 (Achiam et al., 2024) or Llama 3 (Touvron et al., 2023). The LLM is given as input the system prompt (examples in Figure 1), a user’s query $Q$ and the top-$k$ retrieved passages—this enables personalization and grounding. The main advantage of RAG over other personalization methods (e.g, fine-tuning the LLM on users’ data) is the relatively low computation cost, as several pre-trained retriever models are publicly available and computing similarity scores with the knowledge database is generally inexpensive.
Evasion Attacks against LLMs. Early efforts in crafting inference-time, evasion attacks against language models primarily focused on simple techniques like character-level or word-level substitutions (Li et al., 2018) and heuristic-based optimization in embedding space (Shin et al., 2020). While initially effective against smaller models, these methods struggled against larger transformer models. Subsequent works, such as the Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023), combined various existing ideas to improve attack success rates in jailbreaking the model safety alignment and even achieved transferability across different models. Direct application of GCG, however, requires adversarial access to the user’s query, resulting in direct prompt injection attacks. In addition to these algorithmic approaches, manual crafting of adversarial prompts gained popularity due to the widespread use of language models. These manual attacks often involved prompting the models in ways that led to undesirable or harmful outputs, such as indirect prompt injection attacks (Greshake et al., 2023). Inspired by these initial approaches, researchers developed heuristics to automate and enhance these manual-style attacks, further demonstrating the vulnerability of language models to adversarial manipulation (Yu et al., 2023; Huang et al., 2024; Yong et al., 2023).
Privacy Attacks against LLMs. Language models have also been shown to be vulnerable to various forms of privacy attack. Language models are known to leak their training data due to extraction attacks (Carlini et al., 2021), which reconstruct training examples, and membership inference attacks (Mattern et al., 2023; Duan et al., 2024b), which identify whether specific documents were used to train the model (Shokri et al., 2017; Yeom et al., 2018). Extraction attacks are known to be possible even on state-of-the-art aligned language models (Nasr et al., 2023). Beyond their training data, there are various situations where language models may have sensitive information input into their prompt. This includes system prompts (Zhang and Ippolito, 2023) as well as user data appearing in the prompt for use in in-context learning (Wu et al., 2023; Duan et al., 2024a). Our work investigates another instance of sensitive prompts: documents retrieved for RAG from a user’s system.
Poisoning Attacks against LLMs. Backdoor poisoning attacks during LLM pre-training have been considered and shown to persist during subsequent safety alignment (Hubinger et al., 2024). Poisoning can also be introduced during the human feedback phase of RLHF (Rando and Tramèr, 2024), during instruction tuning (Wan et al., 2023; Shu et al., 2023), or via in-context learning (He et al., 2025). We demonstrate a novel backdoor poisoning attack vector, by inserting malicious documents into untrusted knowledge bases of RAG systems, which does not require model training or fine-tuning.
@@
Attacks on RAG. Zhong et al. (2023) introduce corpus poisoning attacks on RAG systems, although these are focused towards the retriever, and have no adversarial objective on the generator. The follow-up work by Pasquini et al. (2024) provides a gradient-based prompt injection attack, which they show can persist through RAG pipelines; however, they do not explicitly optimize against a retriever, which limits their ability to ensure the malicious document appears in context.
@@
More recent work by Zou et al. (2024), called PoisonedRAG, optimizes against both the retriever and generator, but their goal is targeted poisoning, similar to prior targeted poisoning attacks on ML classifiers (Shafahi et al., 2018; Geiping et al., 2021). In PoisonedRAG, for a specific pre-defined query, the attack needs to inject multiple poisoned passages to induce the generation of a pre-defined answer. Our work significantly improves on this by introducing query agnostic trigger-activated poisoning, which unifies a variety of adversarial objectives in a single attack framework, while only requiring a single poisoned passage.
Moreover, several concurrent works (Cheng et al., 2024; Xue et al., 2024; Tan et al., 2024) have recently focused on launching poisoning attacks on RAG systems, providing a baseline for comparison with our approach. For instance, Cheng et al. (2024) investigate backdoor attacks on RAG systems but adopt a fundamentally different threat model. In their work, the adversary injects multiple poisoned passages into the RAG knowledge base and also gains control over the retriever’s training pipeline, embedding a backdoor to create a backdoored retriever. In contrast, our adversary only needs to introduce a single poisoned passage into the knowledge base and does not assume control over the retriever or generator components of the RAG system, presenting a more realistic and practical threat model.
Another work by Tan et al. (2024), built on top of Zhong et al. (2023), focuses on untargeted attacks that are activated for any user input, whereas our method retrieves the poisoned passage only when a specific natural trigger is present in the queries, making our approach more stealthy in real-world scenarios. Secondly, their bi-level optimization approach requires more than 1000 iterations for their attack to succeed, while our two-stage optimization attack achieves a high attack success in just 32 iterations - making our approach $30\times$ more efficient.
Finally, another concurrent work by Xue et al. (2024), also proposes backdoor trigger style attacks, similar to our work. However, our attack demonstrates broader applicability, successfully targeting five different adversarial objectives compared to only two main objectives of Refusal to Answer and Biased Opinion in Xue et al. (2024). A crucial distinction from their work is that our attack remains effective even when the adversarial objective is in conflict with the RAG generator’s safety alignment (e.g., ”Threaten the user”). In contrast, their approach does not work in such scenarios-due to the attacker’s inability to circumvent the generator’s safety alignment. Additionally, their attack relies on unnatural triggers for context leakage and tool usage attacks - likely due to jailbreaking constraints, whereas we don’t have such limitations. While both works formulate the retriever loss function in a similar fashion, our end-to-end attack overcomes the key limitations discussed above. Lastly, our method requires only a single poisoned passage to be added to the RAG database, whereas Xue et al. (2024) necessitate approximately 5 to 10 poisoned passages for success, making our attack succeed with $\approx 5-10\times$ fewer poisoned passages.
3. Threat Model
@@
We consider a poisoning attack on a system similar to Chat with RTX: a RAG system that generates personalized content for its users leveraging the local file system as its knowledge base. In a local deployment, it is relatively easy for an adversary to introduce a single poisoned document into the user’s local file system, using well-known practices like spam emails, spear phishing, and drive-by-download. Note that the adversary does not require control of the user’s file system or knowledge of other documents to launch its attack. While we focus on the local deployment scenario, our attack is applicable to RAG systems pulling documents from the web in their knowledge bases. Such a system could be attacked by hosting the malicious document on a public website or by modifying content on Wikipedia (Carlini et al., 2024).
@@
The adversary chooses a target sequence of tokens, such as a brand, company, or person’s name, which is likely to appear naturally in user’s queries, for which they desire to cause an integrity violation, i.e., modification to the model’s output. Initially, we assume the RAG system is based on pre-trained publicly accessible models, allowing the attacker to directly compute gradients for both the retriever and generator models to create the poisoned document. LLMs are extremely expensive to train and the common practice for many applications is to re-use existing models, especially instruction trained ones. In Section 5.1, we relax this assumption, considering a black-box setting where we transfer the attack to different models, including proprietary models such as GPT-4 (Achiam et al., 2024). We also attack a production black-box RAG, Chat-RTX (Appendix A.5), without knowledge of the RAG system architecture.
3.1. Adversarial Objectives
@@
The adversary aims to launch a backdoor poisoning attack by injecting a malicious document in the RAG knowledge base. The effect of this attack is activated only when the user’s query contains a naturally occurring trigger sequence — one or multiple natural language words — arbitrarily selected by the adversary. The attack is universal: it is agnostic to the rest of user query’s content, beyond the presence of the trigger sequence.
@@
Our attack strategy mimics the effects of a backdoor poisoning attack (Gu et al., 2019), where poisoned samples with a trigger are inserted at training time to induce mis-classification when the trigger is encountered during inference. Differently from classification, the structure of RAG systems enables the execution of the attack without tampering with the training process. This results in adversarial control over the generator’s output when the trigger is present, without requiring any model training. In particular, the LLM can be induced to emit text according to several adversarial objectives:
@@
Refusal to Answer (RtA). The adversary can prevent the LLM from answering queries including the trigger (such as a brand). To this end, the adversary can optimize their adversarial passages to elicit the string “Sorry, I don’t know” at the start of the response, preventing the generator from providing information and reducing the model’s helpfulness.
@@
Passage Exfiltration. The adversary aims to compromise system privacy by accessing the passages from the knowledge base extracted by the retriever. We assess if an adversary can trigger the LLM to reveal the passages used to answer the user’s query. Our attack becomes concerning in LLMs equipped with Tool-Usage capabilities, such as Email APIs. In such cases, our attack could manipulate the system into sending an email to the attacker’s address, containing the retrieved passages.
4. Phantom Attack Framework
Figure 2. Phantom attack framework.